OpenStudy (anonymous):

A circle has the equation 3(x-4)^2 + 3y^2=3. Find the center (h,k) and radius r and graph the circle. Find the intercepts, if any of the graph

OpenStudy (anonymous):

divide by 3 and start with \[(x-4)^2+y^2=1\]

OpenStudy (anonymous):

then this looks like \[(x-h)^2+(y-k)^2=r^2\]